the speed of an airplane in still air is 286 km/h it travels 883 km against the wind and 1119 km with the wind in total time is 13 hr.
Answer by checkley71(8403)

I THINK YOU HAVE SOME OF THE FACTS OR FIGURES WRONG AS THIS PROBLEM HAS NO REASONABLE ANSWER. I GET AN ANSWER OF 286 KM/H FOR THE WIND.
THE EQUATION I CAME UP WITH IS (X^2-22X-75504) WITH FACTORS OF 286 & -264.
-------------------------------------------------------------
A 7 HOUR TRIP MAKES SENSE BUT NOT A 13 HOUR TRIP.
-------------------------------------------------------------
distance=rate*time or distance/rate=time
833/286+1119/286=2.91+3.91=6.82 hours with no wind.
cannot calculate any wind that would extend this time from 6.82 hours to 13 hours.
I suggest you again check your problem --- it doesn't make sense!!!!